OpenStudy (anonymous):

Which of the following points lies on the circle whose center is at the origin and whose radius is 10? (-6, 4) (square root of 10, square root of 10) (6, -8)

OpenStudy (anonymous1):

The equation of the circle whose center is at the origin and whose radius is 10 is: \[x^2+y^2=10^2\] \[x^2+y^2=100\] So, you just have to replace the values of x and y with the abscissa and ordinate of each given point, and see which one gives a correct identity.

OpenStudy (anonymous):

so the second one?

OpenStudy (anonymous1):

Let's check... the point in this case is \[(\sqrt{10},\sqrt{10})\] Let's replace x=square root of 10 and y=square root of 10 in the equation of the circle: \[x^2+y^2=100\] \[\sqrt{10}^2+\sqrt{10}^2=100\] \[10+10=100\] \[20=100\] So, the second one gives a wrong result.

OpenStudy (anonymous1):

The right answer is the third point (6,-8), because \[6^2+(-8)^2=100\] \[36+64=100\] \[100=100\] 100=100 is a true identity. Therefore, the third point is the one we're looking for.